# Temperature dependent energy gap for Yu-Shiba-Rusinov states at the quantum phase transition

### v1

@@ IV.3 Magnetization
In Fig. 4a) we numerically extract the magnetization close to the QPT as a function of $\tilde{J}$ for the same temperatures as in Fig. 3a),c). We find that the magnetization follows a trend very similar to that of the local order parameter $\Delta_{\textbf{x}_{0}}$ around $\tilde{J}_{C}$ in all temperature regimes. Thus, at zero temperature, the magnetization is a step function from zero to $-\frac{1}{2}$ at $\tilde{J}_{C}$. For small but finite temperatures, this step is still present but becomes less pronounced with increasing temperatures. For temperatures $T>T^{*}$, the discontinuity in magnetization disappears and the magnetization instead takes constant value of $\left\langle s_{z}\right\rangle=-\frac{1}{4}$ at $\tilde{J}_{C}$ for all temperatures. In the same way as the local superconducting order parameter is linked to the occupation of the YSR state, so is then the magnetization. At zero temperature, the occupation changes abruptly at $\tilde{J}_{C}$ and hence we have a full-height step function in the magnetization. With increasing temperature, the occupation close to $\tilde{J}_{C}$ is not full for either YSR state and hence the step is smaller. In particular, there is a “rounding" for $\tilde{J}<\tilde{J}_{C}$ since this is where the YSR states are closest to zero energy because of the gap developing on the other side of the phase transition, see Fig. 3a). Finally, at higher temperatures, we reach a regime in which both YSR states cross at zero energy and thus are equally occupied, thus explaining $\left\langle s_{z}\right\rangle=-\frac{1}{4}$.
@@
Based on the results in Fig. 4a), we can view the drop in magnetization as a measure for the QPT, and equivalent for $\tilde{J}_{C}$ (see vertical dotted lines), if the temperature of the system is below $T^{*}$. On the other hand, if there is no clear drop in magnetization at the QPT, the temperature in the system is above $T^{*}$. In Fig. 4b) we show the transition of the system from magnetization 00 to $-0.5$ for different temperatures and magnetic couplings, with the blue shaded area marking the transition region where the magnetization takes intermediate values (using a cut-off of $5\%$ from the limiting 0 and $-0.5$ values). Additionally, we plot the critical coupling $\tilde{J}_{C}$ at different temperatures (solid black line) and also $\tilde{J}_{\pi}$, which is the coupling where $\pi$-shift of the local order parameter sets in (dashed orange line). As discussed above, these latter two curves coincide for low temperatures but start to deviate for $T>T^{*}$, where $T^{*}$ is marked with the horizontal dotted line. By choosing to plot our results as in Fig. 4b), we expose a behavior that is seemingly similar to a generic phase diagram for quantum phase transitions [66] with a wide fan extending from the QPT as the temperature is increased. This similar behavior is intriguing as we only include quantum fluctuations at the mean-field level. However, unlike a typical quantum phase transition phase diagram, the fan region, or intermediate magnetization region for the self-consistent order parameter solution, starts to expands only above temperature $T^{*}$, not at absolute zero. This detail is highlighted by the inset in Fig. 4b) and is due to the magnetization changing discontinuously for temperatures below $T^{*}$. This is in fact a distinguishing feature of the self-consistent order parameter solution over the constant order parameter approximation, as the magnetization only changes discontinuously for $T=0$ for the latter due to it never developing an energy gap for the YSR states.
@@ IV.4 Parameter dependence of gap
First of all, for line b) in Fig. 5 we introduce a small, but finite onsite potential of $\tilde{\varepsilon}_{i}=-0.05\,t$ to the impurity site, which is $10\%$ of the chemical potential $\mu$. Such local onsite potential can easy be present for an impurity and should generally be included. We find that the introduction of this local perturbation in the onsite energy at the impurity site has the overall effect of slightly increased gap for most temperatures. Importantly, the energies of the YSR states are now also finite at zero temperature. Generally speaking this is an expected development, as we should expect having YSR states touch zero energy before suddenly jumping to finite energies, as in Fig. 3a), should be a reasonably unstable situation, energy-wise. Here, we see that just perturbing the impurity with a slight shift in the onsite potential lifts the zero-energy states and thus generates a gap also at zero temperature.
@@
We further test our results against changes in the overall chemical potential of the substrate with typical results given by line c) in Fig. 5, using $\mu=3.618\,t$, where the Fermi surface becomes circular in the normal state [31]. Notably, here we find a substantially increased YSR energy gap, around a factor of three times larger by just changing the Fermi surface. The range of temperatures in terms of $T_{C}$, at which the energy gap can be observed at the QPT is further increased by a similar degree. We also observe a finite energy of the YSR at zero temperature, which we here attribute to an increased particle-hole asymmetry of the normal state, not just at the impurity site. We, therefore, conclude that considering a self-consistent superconducting order parameter will in general result in an energy gap for the YSR states at all temperatures $T<T^{*}$, unless a very high degree of particle-hole symmetry is present in the substrate and the impurity.
@@
Finally, we consider changing the superconducting order parameter value, or equivalently the potential $V_{sc}$. We here note that calculations with order parameter sizes closer to existing conventional $s$-wave superconductors are hard to achieve numerically, as such small superconducting gaps demand the modeling of very large system sizes, in order to have a reasonable large number of states at the Fermi surface to condense into the superconducting state. If this is not fulfilled, the results become unstable and inconsistent. We still try to estimate the effects of smaller superconducting gaps by lowering the order parameters to 75% and 50% of the original values for lines d) and e) in Fig. 3d), respectively. The trends that we extract is that the energy gap is decreased with a decreasing order parameter strengths. We further note that, even for this limited range of order parameters, this trend seems to be non-linear. In summary, we find that the YSR energy gap at the QPT, and the temperature range over which it can be observed, is strongly dependent on physical parameters non-related to the magnetic moment of the impurity itself, such as the chemical potential, the onsite impurity energy, and the superconducting bulk order parameter.
@@ V Concluding remarks
We investigate the YSR state spectrum for a magnetic impurity absorbed on a conventional spin-singlet $s$-wave superconductor, with a focus on the regime around the critical coupling where the system undergoes a quantum phase transition (QPT) from a weakly coupled impurity spin to a screened spin state. According to tentative experimental observations [20], the YSR states may remain at finite energy at the QPT, an effect that is not predicted by the original YSR theory. In this work, we identify the only two mechanisms that can cause cause finite energy YSR states at the QPT: by introducing additional interactions to the model description or, by abrupt changes in the system triggered by the QPT.
@@
We showed first that adding additional interactions on the impurity itself or the bond do not cause a finite YSR energy at the QPT. The only exception is if the substrate lacks inversion symmetry, for example, due to strong disorder, or if the impurity coupling is not fully localized in the substrate while also breaking the inversion symmetry. Thus, the only opportunity to generate a gap in the YSR state spectrum is following the second mechanism, and, in particular, explore if the local abrupt changes in the superconducting order parameter of the substrate at the QPT can induces a gap. Such changes are known to exist [22, 23, 24, 30, 25, 26, 31, 32, 33] and our results show that they also robustly induce a gap in the YSR spectrum below a certain temperature, $T^{*}$.
@@
To summarize the reason and behavior of the gap, at finite temperatures below $T^{*}$, there is a discontinuity between the values of the local order parameter for magnetic couplings below and above the critical magnetic coupling at the QPT. This not only leads to an order parameter with opposite sign at the impurity site, a so-called $\pi$-shift reported previously [22, 23, 24, 30, 25, 26, 31, 32, 33], but also to a shift in the spectrum of the YSR states. This spectral shift causes the YSR states to “miss" zero energy and, therefore, remain at finite energies even at the critical coupling strength. Overall, this gives rise to a gap, and often also to a discontinuous jump in the YSR energies.
@@
We further find that the temperature $T^{*}$ is highly dependent on the model parameters. As some of these parameters are challenging to determine or vary in experimental setups, quantitative predictions about the effects of local suppression of the order parameter and thus the size of the energy gap are difficult to make, given our general model. Further, given the strong dependence on the chemical potential, we expect that the results also strongly depend on the underlying band structure of the normal state at the Fermi surface. Still, the qualitative behavior of the temperature-dependent energy gap developed for the YSR states at the QPT holds for a wide range of possible parameters.
@@ Appendix A Symmetry forbidden gap opening terms
We start by considering a coupling that acts on the particle-particle (hole-hole) part of the Nambu space, i.e. proportional to $\tau_{3}$, but now with a spin-dependence in constrast to the $\varepsilon_{i}$ term in Eq. (1). Here we find that an additional coupling of the form $\tilde{V}_{is}=\alpha_{3}\leavevmode\nobreak\ \sigma_{1,2}\otimes\tau_{3}$ produces an effective interaction between the YSR states at the QPT. To derive analytical results, we keep $\alpha_{3}$ as a generic complex valued scalar and assume that it is much smaller than the superconducting parameter ($\alpha_{3}\ll\Delta$) and that the energies of the resulting YSR states are close to zero energy ($E\ll\Delta$). We then find, for an arbitrary linear combination of the $\sigma_{1}$ and $\sigma_{2}$ spin orientations, in the interaction the following YSR energies
@@ Appendix B Free energy across QPT
Figure 6: a) Free energy contributions $F_{\tilde{\mathcal{H}}}$ and $F_{\Delta}$ at zero temperature as a function of $\tilde{J}$ and normalized by the number of sites $N$. Dashed vertical line denotes $\tilde{J}_{C}$. b) Spatial behavior of the local order parameter along the [11] direction. Two values, $\tilde{J}_{1}$ and $\tilde{J}_{2}$, are chosen to be just below and above $\tilde{J}_{C}$ within the numerical resolution of $\tilde{J}$. Solid black line indicate bulk $\Delta$ (for $\tilde{J}=0)$. Green line shows the difference in $\Delta_{\textbf{i}}$ between these two cases.
@@
where the first contribution $F_{\tilde{\mathcal{H}}}$ comes from the energy spectrum of the Hamiltonian $\tilde{\mathcal{H}}$ given the sum of the Hamitonians in Eq. (11) and Eq. (9), and with $f_{0}$ being the zero temperature Dirac-Fermi function. The second part, $F_{\Delta}$, accounts for additional energy contribution of the superconducting order parameter due to its mean-field decomposition. For both of these contributions, we find a jump at the QPT at $\tilde{J}_{C}$, significantly lowering the energy of the ground state for $\tilde{J}>\tilde{J}_{C}$, see Fig. 6a). The drop in $F_{\tilde{\mathcal{H}}}$ is largest and is a direct consequence of jump of the YSR energies at $\tilde{J}_{C}$. The drop in $F_{\Delta}$ is smaller, where a drop, instead of a raise, may at first seem surprising since $\Delta$ is drastically suppressed at the impurity site. To understand this we plot in 6b) the local order parameter extracted along the [11] direction, cutting through the impurity. There is a strong local suppression of the order parameter together with a $\pi$-shift at the impurity site, which itself yields an increase in the free energy for $\tilde{J}>\tilde{J}_{C}$ compared to $\tilde{J}<\tilde{J}_{C}$. However, we find that the suppression of the local order parameter becomes more localized and the order parameter is even slightly enhanced in the vicinity of the impurity for magnetic coupling is above $\tilde{J}_{C}$. Similar “ripples" have been reported before for magnetic impurities [28, 22, 23, 24]. Their exact location is due to the YSR state localization being dependent on $v_{F}$. Overall, we find that these effects actually outweigh the local suppression of $\Delta_{\textbf{x}_{0}}$. It is most pronounced along the [11] direction of the lattice but exists along other directions as well. This directional dependence is due to the normal-state Fermi surface. As expected, for a fully circular Fermi surface, we find a circular suppression/enhancement of $\Delta_{\textbf{i}}$. We observed a similar behavior for the free energy for finite temperatures, while the additional free energy contribution of the entropy is negligible.

### v2

@@ IV.3 Magnetization
In Fig. 4a) we numerically extract the magnetization close to the QPT as a function of $\tilde{J}$ for the same temperatures as in Fig. 3a),c). We find that the magnetization follows a trend very similar to that of the local order parameter $\Delta_{\textbf{x}_{0}}$ around $\tilde{J}_{C}$ in all temperature regimes. Thus, at zero temperature, the magnetization is a step function from zero to $\frac{1}{2}$ at $\tilde{J}_{C}$. For small but finite temperatures, this step is still present but becomes less pronounced with increasing temperatures. For temperatures $T>T^{*}$, the discontinuity in magnetization disappears and the magnetization instead takes constant value of $\left\langle s_{z}\right\rangle=-\frac{1}{4}$ at $\tilde{J}_{C}$ for all temperatures. In the same way as the local superconducting order parameter is linked to the occupation of the YSR state, so is then the magnetization. At zero temperature, the occupation changes abruptly at $\tilde{J}_{C}$ and hence we have a full-height step function in the magnetization. With increasing temperature, the occupation close to $\tilde{J}_{C}$ is not full for either YSR state and hence the step is smaller. In particular, there is a “rounding" for $\tilde{J}<\tilde{J}_{C}$ since this is where the YSR states are closest to zero energy because of the gap developing on the other side of the phase transition, see Fig. 3a). Finally, at higher temperatures, we reach a regime in which both YSR states cross at zero energy and thus are equally occupied, thus explaining $\left\langle s_{z}\right\rangle=-\frac{1}{4}$.
@@
Based on the results in Fig. 4a), we can view the drop in magnetization as a measure for the QPT, and equivalent for $\tilde{J}_{C}$ (see vertical dotted lines), if the temperature of the system is below $T^{*}$. On the other hand, if there is no clear drop in magnetization at the QPT, the temperature in the system is above $T^{*}$. In Fig. 4b) we show the transition of the system from magnetization 0 to 0.5-0.5 for different temperatures and magnetic couplings, with the blue shaded area marking the transition region where the magnetization takes intermediate values (using a cut-off of $5\%$ from the limiting 0 and 0.5-0.5 values). Additionally, we plot the critical coupling $\tilde{J}_{C}$ at different temperatures (solid black line) and also $\tilde{J}_{\pi}$, which is the coupling where $\pi$-shift of the local order parameter sets in (dashed orange line). As discussed above, these latter two curves coincide for low temperatures but start to deviate for $T>T^{*}$, where $T^{*}$ is marked with the horizontal dotted line. By choosing to plot our results as in Fig. 4b), we expose a behavior that is seemingly similar to a generic phase diagram for quantum phase transitions [68] with a wide fan extending from the QPT as the temperature is increased. This similar behavior is intriguing as we only include quantum fluctuations at the mean-field level. However, unlike a typical quantum phase transition phase diagram, the fan region, or intermediate magnetization region for the self-consistent order parameter solution, starts to expands only above temperature $T^{*}$, not at absolute zero. This detail is highlighted by the inset in Fig. 4b) and is due to the magnetization changing discontinuously for temperatures below $T^{*}$. This is in fact a distinguishing feature of the self-consistent order parameter solution over the constant order parameter approximation, as the magnetization only changes discontinuously for $T=0$ for the latter due to it never developing an energy gap for the YSR states.
@@ IV.4 Parameter dependence of gap
First of all, for line b) in Fig. 5 we introduce a small, but finite onsite potential of $\tilde{\varepsilon}_{i}=-0.05\,t$ to the impurity site, which is $10\%$ of the chemical potential $\mu$. Such local onsite potential can easy be present for an impurity and should generally be included. We find that the introduction of this local perturbation in the onsite energy at the impurity site has the overall effect of slightly increased gap for most temperatures. Importantly, the energies of the YSR states are now also finite at zero temperature. Generally speaking this is an expected development, as we should expect having YSR states touch zero energy before suddenly jumping to finite energies, as in Fig. 3a), to be an unstable situation, energy-wise. Here, we see that perturbing the impurity with just a slight shift in the onsite potential lifts the zero-energy states and thus generates a gap also at zero temperature.
@@
We further test our results against changes in the overall chemical potential of the substrate with typical results given by line c) in Fig. 5, using $\mu=3.618\,t$, where the Fermi surface becomes circular in the normal state [32]. Notably, here we find a substantially increased YSR energy gap, around a factor of three times larger by just changing the Fermi surface. The range of temperatures in terms of $T_{C}$, at which the energy gap can be observed at the QPT, is further increased by a similar degree. We also observe a finite energy of the YSR states at zero temperature, which we here attribute to an increased particle-hole asymmetry of the normal state, not just at the impurity site. This can be explained by the fact that particle-hole asymmetry lowers the energy degeneracy of the YSR states at zero energy and thus better facilitates lowering of the total free energy with a gap opening. In fact, we generally find a stronger suppression of the superconducting order parameter with increased particle-hole asymmetry of the normal state, directly associated with the increased YSR gap. We, therefore, conclude that considering a self-consistent superconducting order parameter will in general result in an energy gap for the YSR states at all temperatures $T<T^{*}$, unless a very high degree of particle-hole symmetry is present in the substrate and the impurity. We also find that the overall shape of the Fermi surface may influence the size of the gap. This is expected since the Fermi surface is known to influence the properties of the YSR states [23, 24, 22, 31, 32, 53].
@@
Finally, we consider changing the superconducting order parameter value, or equivalently the potential $V_{sc}$. We here note that calculations with order parameter sizes closer to existing conventional $s$-wave superconductors are hard to achieve numerically, as such small superconducting gaps demand the modeling of very large system sizes, in order to have a reasonable large number of states at the Fermi surface to condense into the superconducting state. If this is not fulfilled, the results become unstable and inconsistent. We still try to estimate the effects of smaller superconducting gaps by lowering the order parameters to 75% and 50% of the original values for lines d) and e) in Fig. 3d), respectively. The trends that we extract is that the energy gap is decreased with a decreasing order parameter strengths. We further note that, even for this limited range of order parameters, this trend seems to be non-linear.
In summary, we find that the YSR energy gap at the QPT, and the temperature range over which it can be observed, is strongly dependent on physical parameters non-related to the magnetic moment of the impurity itself, such as the chemical potential, the onsite impurity energy, the Fermi surface shape, and the superconducting bulk order parameter. Thus, we conclude that the exact value of the energy gap between the YSR states, as well as the transition temperature $T^{*}$ is highly dependent on the band structure of the substrate and the exact value of the bulk order parameter. Therefore, our results are qualitative. For quantitative predictions, the band structure of the material and impurity involved should be taken into account. In particular, particle-hole asymmetry in the normal state has a tendency to increase the YSR energy gap.
@@ V Concluding remarks
We investigate the YSR state spectrum for a magnetic impurity absorbed on a conventional spin-singlet $s$-wave superconductor, with a focus on the regime around the critical coupling where the system undergoes a quantum phase transition (QPT) from a weakly coupled impurity spin to a screened spin state. According to tentative experimental observations [20], the YSR states may remain at finite energy at the QPT, an effect that is not predicted by the original YSR theory. In this work, we identify the only two mechanisms that may cause cause finite energy YSR states at the QPT: by introducing additional interactions to the model description or, by abrupt changes in the system, triggered by the QPT.
@@
We first showed that adding additional interactions on the impurity itself or its bond to the substrate do not cause a finite YSR energy at the QPT. The only exception is if the substrate lacks inversion symmetry, for example, due to strong disorder, or if the impurity coupling is not fully localized in the substrate and at the same time also breaking inversion symmetry. In typical experimental configurations, including the one in which finite YSR state energies has been documented [20], inversion symmetry is seemingly intact and thus additional interactions are not causing the YSR energy gap. Thus, we conclude that the YSR gap is much more likely generated by the second mechanism, which evokes an abrupt change to the system triggered at the QPT. The only parameter to undergo such an abrupt change is the superconducting order parameter, which self-consistently responds to the YSR states and is thus particularly sensitive to the crossing of YSR states at the QPT. Such abrupt changes of the superconducting order parameter at the QPT are already known to exist, often even resulting in a local $\pi$-shift at the impurity site [22, 23, 24, 31, 48, 26, 32, 33, 34]. Our results show that the self-consistently heavily suppressed superconducting order parameter at the impurity also robustly induce a gap in the YSR spectrum below a certain temperature, $T^{*}$.
@@
To summarize the reason and behavior of the induced gap: at finite temperatures below the temperature $T^{*}$, there is a discontinuity between the values of the local order parameter for magnetic couplings below and above the critical magnetic coupling at the QPT. This not only leads to an order parameter with opposite sign at the impurity site, a so-called $\pi$-shift reported previously [22, 23, 24, 31, 48, 25, 26, 32, 33, 34], but also to a shift in the spectrum of the YSR states. This spectral shift causes the YSR states to “miss" zero energy and, therefore, remain at finite energies even at the critical coupling strength. Overall, this gives rise to a gap, and often also to a discontinuous jump in the YSR energies.
@@
We further find that the temperature $T^{*}$ is highly dependent on the model parameters. As some of these parameters are challenging to determine or vary in experimental setups, quantitative predictions about the effects of local suppression of the order parameter and thus the size of the energy gap are difficult to make, given our general model. Further, given the strong dependence on the chemical potential, we expect that the results also strongly depend on the underlying band structure of the normal state at the Fermi surface. Still, the qualitative behavior of a strongly temperature-dependent energy gap for the YSR states at the QPT holds for the full range of parameters we explore.
@@
Lastly, we would like to point out that our model description includes fluctuations of the magnetic impurity only at the mean field level. Still, our description yield results that are in good agreement with experimental findings [15, 20]. Further including quantum fluctuations of the impurity might modify the results presented in this work. Therefore, a possible future extension of this work could include those effects, for example via NRG calculations.
@@ Appendix A Symmetry forbidden gap opening terms
We start by considering a coupling that acts on the particle-particle (hole-hole) part of the Nambu space, i.e. proportional to $\tau_{3}$, but now with a spin-dependence in constrast to the $\varepsilon_{i}$ term in Eq. (1). Here we find that an additional coupling of the form $\tilde{V}_{is}=\alpha_{3}\penalty 10000\ \sigma_{1,2}\otimes\tau_{3}$ produces an effective interaction between the YSR states at the QPT. To derive analytical results, we keep $\alpha_{3}$ as a generic complex valued scalar and assume that it is much smaller than the superconducting parameter ($\alpha_{3}\ll\Delta$) and that the energies of the resulting YSR states are close to zero energy ($E\ll\Delta$). We then find, for an arbitrary linear combination of the $\sigma_{1}$ and $\sigma_{2}$ spin orientations, in the interaction the following YSR energies
@@ Appendix B Free energy across QPT
Figure 6: a) Free energy contributions $F_{\tilde{\mathcal{H}}}$ and $F_{\Delta}$ at zero temperature as a function of $\tilde{J}$ and normalized by the number of sites NN. Dashed vertical line denotes $\tilde{J}_{C}$. b) Spatial behavior of the local order parameter along the [11] direction. Two values, $\tilde{J}_{1}$ and $\tilde{J}_{2}$, are chosen to be just below and above $\tilde{J}_{C}$ within the numerical resolution of $\tilde{J}$. Solid black line indicate bulk $\Delta$ (for $\tilde{J}=0)$. Green line shows the difference in $\Delta_{\textbf{i}}$ between these two cases.
@@
where the first contribution $F_{\tilde{\mathcal{H}}}$ comes from the energy spectrum of the Hamiltonian $\tilde{\mathcal{H}}$ given the sum of the Hamiltonians in Eq. (11) and Eq. (9), and with $f_{0}$ being the zero temperature Dirac-Fermi function. The second part, $F_{\Delta}$, accounts for additional energy contribution of the superconducting order parameter due to its mean-field decomposition. For both of these contributions, we find a jump at the QPT at $\tilde{J}_{C}$, significantly lowering the energy of the ground state for $\tilde{J}>\tilde{J}_{C}$, see Fig. 6a). The drop in $F_{\tilde{\mathcal{H}}}$ is largest and is a direct consequence of jump of the YSR energies at $\tilde{J}_{C}$. The drop in $F_{\Delta}$ is smaller, where a drop, instead of a raise, may at first seem surprising since $\Delta$ is drastically suppressed at the impurity site. To understand this we plot in 6b) the local order parameter extracted along the [11] direction, cutting through the impurity. There is a strong local suppression of the order parameter together with a $\pi$-shift at the impurity site, which itself yields an increase in the free energy for $\tilde{J}>\tilde{J}_{C}$ compared to $\tilde{J}<\tilde{J}_{C}$. However, we find that the suppression of the local order parameter becomes more localized and the order parameter is even slightly enhanced in the vicinity of the impurity for magnetic coupling is above $\tilde{J}_{C}$. Similar “ripples" have been reported before for magnetic impurities [29, 22, 23, 24]. Their exact location is due to the YSR state localization being dependent on $v_{F}$. Overall, we find that these effects actually outweigh the local suppression of $\Delta_{\textbf{x}_{0}}$. It is most pronounced along the [11] direction of the lattice but exists along other directions as well. This directional dependence is due to the normal-state Fermi surface. As expected, for a fully circular Fermi surface, we find a circular suppression/enhancement of $\Delta_{\textbf{i}}$. We observed a similar behavior for the free energy for finite temperatures, while the additional free energy contribution of the entropy is negligible.